1. XYZ par value and market value of its shares are 10$ and 20$ respectively. XYZ issued a stock dividend. Assume XYZ issued 250,000 shares.If the stock dividend amount to 50,000 shares. What is the stock dividend percentage?

2. Target shares have a par value of 5$ per share. Target repurchased 10,000 of its shares at 8$ on September 14, 2018. On November 3, 2019 they sold 5,000 of the same shares for 9$ per share. How much is the additional paid in capital? (treasury stock)

3. 2017 stockholders equity:500,000

2018 stockholders equity:600,000

Net Sales= 200,000

cost of goods sold= 100,000

Operating expenses= 50,000

Income tax Expenses= 20% of the income after all other expenses

What is the return on Equity?
Answer #1

1) Stock dividend percentage = 50000/250000 = 20%

2) Additional paid in capital (treasury stock) = (9-8)*5000 = 5000

3) Net income after tax = 40000

Average equity = 1100000/2 = 550000

return on equity = 40000/550000 = 7.27%